# SpikingResformer: Bridging ResNet and Vision Transformer in Spiking Neural Networks

### v1

@@ 2 Related Work
Spiking Convolutional Neural Networks. Spiking convolutional neural networks (SCNNs) have been extensively developed due to the remarkable success of surrogate gradient learning [42, 30] and are widely used in handling challenging vision tasks, including object recognition [1, 38, 16, 19], detection [15, 35], and segmentation [17, 31]. To improve the performance of SCNNs on these challenging tasks, researchers have dedicated great efforts to exploring training methods [20, 43, 5, 39, 10, 28] and ANN-to-SNN conversion techniques [34, 12, 11, 21, 23, 14, 40, 13, 2, 4]. Moreover, many deep spiking convolutional architectures [34, 20, 43, 12, 7] have been proposed to achieve high performance. The success of SpikingVGG [34, 20] demonstrates that SCNNs can achieve comparable performance to ANNs in recognition tasks. SpikingResNet [43, 12] further explores SCNNs with residual structure and achieves deeper SCNN with ResNet-based architecture. Moreover, SEW ResNet [7] meticulously analyzes the identity mapping in directly-trained spiking residual networks and successfully train a 152-layer SCNN directly. These architectures leverage large-scale SNNs with numerous layers and demonstrate superior performance on various tasks.
@@ 5.2 Spiking Resformer Block
Spike-driven Transformer-8-768 4 66.34 -/- 6.09/- ${}^{\dagger}$
SpikingResformer (Ours) Direct Training SpikingResformer-Ti 4 11.14 ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
SpikingResformer-S 4 17.76 ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
SpikingResformer-M 4 35.52 ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
SpikingResformer-L 4 60.38 ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (62176003, 62088102) and by Beijing Nova Program (20230484362).
@@ E Experiment Details
SpikingResformer-Ti 4 11.14 ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
SpikingResformer-S 4 17.76 ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
SpikingResformer-M 4 35.52 ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
SpikingResformer-L 4 60.38 ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$

### v2

@@ 2 Related Work
Spiking Convolutional Neural Networks. Spiking convolutional neural networks (SCNNs) have been extensively developed due to the remarkable success of surrogate gradient learning [43, 31] and are widely used in handling challenging vision tasks, including object recognition [1, 20], detection [17, 36], and segmentation [18, 32]. To improve the performance of SCNNs on these challenging tasks, researchers have dedicated great efforts to exploring training methods [21, 6, 39, 12, 29, 47, 48, 41] and ANN-to-SNN conversion techniques [35, 14, 13, 22, 24, 16, 40, 15, 3, 2, 7, 5]. Moreover, many deep spiking convolutional architectures [35, 21, 44, 14, 9] have been proposed to achieve high performance. The success of SpikingVGG [35, 21] demonstrates that SCNNs can achieve comparable performance to ANNs in recognition tasks. SpikingResNet [44, 14] further explores SCNNs with residual structure and achieves deeper SCNN with ResNet-based architecture. Moreover, SEW ResNet [9] meticulously analyzes the identity mapping in directly-trained spiking residual networks and successfully trains a 152-layer SCNN directly. These architectures leverage large-scale SNNs with numerous layers and demonstrate superior performance on various tasks.
@@ 5.2 Spiking Resformer Block
Spike-driven Transformer-8-768 4 66.34 -/- 6.09/- 76.32/77.07
SpikingResformer (Ours) Direct Training SpikingResformer-Ti 4 11.14 2.73/4.71 2.46/4.24 74.34/75.57
SpikingResformer-S 4 17.76 3.74/6.40 3.37/5.76 75.95/76.90
SpikingResformer-M 4 35.52 6.07/10.24 5.46/9.22 77.24/78.06
SpikingResformer-L 4 60.38 9.74/16.40 8.76/14.76 78.77/79.40
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (62176003, 62088102) and by the Beijing Nova Program (20230484362).
@@ E Experiment Details
SpikingResformer-Ti 4 11.14 2.73/4.71 2.46/4.24 74.34/75.57
SpikingResformer-S 4 17.76 3.74/6.40 3.37/5.76 75.95/76.90
SpikingResformer-M 4 35.52 6.07/10.24 5.46/9.22 77.24/78.06
SpikingResformer-L 4 60.38 9.74/16.40 8.76/14.76 78.77/79.40